# arXiv:2603.01390v1[math.RT] 02 Mar 2026

### v1

@@
Duflo–Serganova functors play a central role in the representation theory of Lie superalgebras and have been studied extensively, especially in the finite-dimensional setting. On the other hand, in the general linear case, Brundan–Goodwin observed that certain distinguished parabolic induction functors are particularly compatible with the principal Whittaker coinvariants functor $H_{0}$, and these functors are crucial in their theory.
In this paper we show that, for certain rank-one DS functors, the image under DS can be computed explicitly for a class of infinite-dimensional modules which can be realized as images of these parabolic induction functors. In particular, for $\mathfrak{b}$-Verma supermodules attached to a suitable class of Borel subalgebras $\mathfrak{b}$ and a family of modules whose $H_{0}$-images are tensor products of one-dimensional evaluation modules for the super Yangian, we obtain an elegant explicit formula.
@@ 1 Introduction
Highest weight representations, together with BGG category $\mathcal{O}$, are among the most fundamental and central objects in the representation theory of semisimple Lie algebras. In this area, endofunctors attached to roots—such as translation functors—have played a crucial role for decades [21]. Representation theory of basic classical Lie superalgebras, in particular for the general linear Lie superalgebra $\mathfrak{gl}(m|n)$ [6], provides a natural extension of the classical theory and exhibits phenomena absent in the Lie algebra case, such as the fact that Borel subalgebras need not be conjugate. Nevertheless, many of the classical constructions still make sense. For instance, twisting functors associated with even roots have been studied extensively [10]. It is therefore natural to ask whether there exist similarly well-behaved functors associated with odd roots, a question which is expected to reveal phenomena specific to the super setting.
@@
The Duflo–Serganova functor [12] is widely regarded as the canonical example—and perhaps the only currently known example—of a “good” functor attached to an odd root. For $\mathfrak{gl}(m|n)$ it maps representations to those of a smaller Lie superalgebra $\mathfrak{gl}(m-r|n-r)$, thereby reducing complexity. This functor is symmetric monoidal and has left a deep imprint on the theory of finite-dimensional representations; we do not attempt to survey this vast literature here, and instead refer the reader to the survey [14]. For the purposes of this paper, we single out one striking instance: Heidersdorf–Weissauer [16] determine the images of finite-dimensional irreducible $\mathfrak{gl}(n|n)$-modules under Duflo–Serganova functors in terms of Khovanov arc diagrams.
@@
In contrast, much less is known on the infinite-dimensional side. In [11], Coulembier and Serganova study which Verma modules attached to the distinguished Borel are annihilated by Duflo–Serganova functors, and apply this to homological-algebraic questions (these results have been generalized to arbitral Borel subalgebras in part by the author in [19]). Hoyt–Penkov–Serganova [20] show that Duflo–Serganova functors preserve category $\mathcal{O}$ and formulate further conjectures on their behavior. Taken together, these developments suggest that it is natural to investigate the behavior of Duflo–Serganova functors on category $\mathcal{O}$ beyond the finite-dimensional setting. Nevertheless, to the best of the author’s knowledge, concrete nontrivial explicit computations for infinite-dimensional highest weight modules have not appeared. The aim of this paper is to exhibit a class of infinite-dimensional highest weight representations of $\mathfrak{gl}(n|n)$ for which explicit computations with a rank-one Duflo–Serganova functor become possible.
@@
Hypercubic decomposition of Verma modules. A first candidate would be the $\mathfrak{b}$-Verma modules for a Borel $\mathfrak{b}$. However, in view of the fact that DS functors are middle exact, it is natural to decompose $\mathfrak{b}$-Verma modules along the directions of odd $\mathfrak{b}$-simple roots. In [19], the author formalized this idea by realizing, inside $\mathcal{O}_{\mathfrak{gl}(n|n)}$, certain abelian subcategories which are equivalent to blocks of $\mathcal{O}_{\mathfrak{gl}(1|1)^{\oplus n}}$ and which contain the relevant modules. Variants of this idea already appear in the literature, for instance in work of Cheng–Lam–Wang [8] and Serganova [24]. Pursuing this direction leads to a remarkable family of hypercube Borels, for which some of the associated $\mathfrak{b}$-Verma modules can be realized inside an abelian subcategory equivalent to the maximal atypical block of $\mathcal{O}_{\mathfrak{gl}(1|1)^{\oplus n}}$. In this sense, this situation provides the finest possible decomposition of $\mathfrak{b}$-Verma modules.
@@
For example, already for $\mathfrak{gl}(3|3)$ one can visualize the collection of Borel subalgebras with fixed even part as a finite graph whose vertices are Borels and whose edges correspond to odd reflections. In this picture there is a distinguished “cube” in the middle, and the remarkable family of hypecube Borels refers to the Borel subalgebras corresponding to the vertices of this central cube. Note that here the distinguished Borel corresponds to $()$.
[…]
Such Borels have appeared in several places; for example, they are used essentially in Cheng–Lam’s proof of super duality [9].
Parabolic induction and Whittaker coinvariant functor.
@@
Studying the decomposition of Verma modules for hypercube Borels is equivalent to considering the parabolic induction functor of Brundan–Goodwin [4], which sends $\mathfrak{gl}(1|1)^{\oplus n}$-modules to $\mathfrak{gl}(n|n)$-modules. The relevance of these parabolic inductions is explained by the principal Whittaker coinvariant functor $H_{0}$, which sends objects of category $\mathcal{O}$ to modules over the (nontrivial!) principal $W$-superalgebra. Generalizing [5], Brown–Brundan–Goodwin [3] show that the principal $W$-superalgebra can be realized as a quotient of the super Yangian $Y(\mathfrak{gl}(1|1))$.
Brundan–Goodwin study $H_{0}$ by exploiting the fact that the above parabolic induction images are sent, under $H_{0}$, to tesor product of evaluation modules for the super Yangian. In particular, they show that Verma modules for the distinguished Borel are mapped to Verma modules on the $W$-superalgebra side, defined via the Drinfeld presentation of super Yangians, and that the Whittaker coinvariant functor $H_{0}$ provides a realization of the Soergel functor. Thus these parabolic induction functors, and their images, play an essential role in the study of category $\mathcal{O}$; we return to this point in §3.
What we prove. In §5 we show that, for a certain rank-one Duflo–Serganova functor, the image under DS can be computed explicitly on a class of modules containing the images of the parabolic induction functors considered by Brundan–Goodwin. In particular, in a setting which includes the case where $\mathfrak{b}$ is a hypercube Borel and $e_{\alpha}=e_{1,n+1}$, we prove the following conjecture.
@@
In §6 we verify this conjecture in the case of $\mathfrak{gl}(2|2)$ by direct calculations using Hinich’s lemma.
As suggested by the conjecture, our computations for $\mathfrak{b}$-Verma modules are not expected to depend essentially on our particular choice of Borel subalgebras. However, an important feature of our hypercube Borels is the following result.
@@
$\operatorname{DS}_{e_{1,n+1}}\operatorname{BG}_{n|n}(\lambda)\ \cong\ \operatorname{BG}_{n-1|n-1}\bigl(\operatorname{pr}_{\alpha}(\lambda)\bigr),$
and moreover $H_{0}\operatorname{BG}_{n-1|n-1}\bigl(\operatorname{pr}_{\alpha}(\lambda)\bigr)$ is also one-dimensional.
Remarks. Although we expect that our results can be generalized to the $\mathfrak{gl}(m|n)$ setting, we believe that, for the development of this area, it is important to focus on $\mathfrak{gl}(n|n)$. Working with $\mathfrak{gl}(n|n)$ brings a certain simplicity and beauty. For instance, in the $\mathfrak{gl}(n|n)$ case one does not need to consider shifts of the super Yangian [3]. Also, a number of existing works suggest that phenomena for $\mathfrak{gl}(m|n)$ often reduce to, or are controlled by, the $\mathfrak{gl}(n|n)$ case; see for example [16, 17, 25].
The main objects of study in this paper, namely the modules $\operatorname{BG}(\lambda)$, can be realized as images of homomorphisms between Verma modules attached to Borel subalgebras lying at opposite vertices of the central hypercube. On the other hand, in [18] the images of homomorphisms between Verma modules attached to the distinguished and anti-distinguished Borels were studied, and were used in a certain generalization of a classical BGG-type realization. In a sense, these two constructions may be viewed as lying at opposite extremes.
Finally, we note that, from the perspective of Nichols algebras of diagonal type and their classification via Weyl groupoids (a.k.a. generalized root systems), changes of Borel are of independent interest (see [2, 1, 15]. We believe that the results of this paper (Theorems 3.13 and 5.9 provide further evidence for the significance of such changes of Borel.
@@ 1.1 Acknowledgments
We define the category of modules with an even Verma flag, $\mathcal{F}\!\Delta_{\bar{0}}\subset\mathcal{O}_{\bar{0}}$.
@@
    $L^{()}(\lambda)$ injectively maps to some $M$ s.t. $\operatorname{Res}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}M\in\mathcal{F}\!\Delta_{\bar{0}}.$
@@ 3 Brundan-Goodwin functors
Under the principal Whittaker coinvariants functor $H_{0}$, the module $\operatorname{BG}(\lambda)$ is sent to the corresponding irreducible evaluation module for the (super) Yangian. Moreover,
@@
    $\operatorname{BG}(\lambda)\ \cong\ \operatorname{Im}\!\left(M^{\mathfrak{b}_{\mathbf{o}}}(\lambda)\xrightarrow{\ \neq 0\ }M^{\mathfrak{b}_{\mathbf{i}}}(\lambda)\right).$
@@ 4 Duflo-Serganova functors
Forx in X , $M\in\mathfrak{g}\text{-sMod}$, we define a supervector space $DS_{x}M$ by:
@@
naturally inherits the structure of a basic Lie superalgebra. Consequently, there is a natural symmetric monoidal $k$-linear functor
@@ 5 Main results
Let $\mathfrak{b}_{1|1}=(\beta_{n})\in L(1,1)$ and $\mathfrak{b}_{n-1|n-1}=(\beta_{1},\dots,\beta_{n-1},\beta_{n}=1)\in L(n-1,n-1)$. Define
@@
For $\lambda\in\Lambda_{n|n}$ we write $\lambda|_{n-1|n-1}\in\Lambda_{n-1|n-1}$ and $\lambda|_{1|1}\in\Lambda_{1|1}$ for the restrictions corresponding to the Levi embedding
@@
$\lambda|_{n-1|n-1}\in\Lambda^{\operatorname{BG}}_{n-1|n-1}\qquad\text{(resp.\ }\lambda|_{n-1|n-1}\in\Lambda^{\operatorname{maBG}}_{n-1|n-1}\text{)}.$
@@
Let $\mathfrak{b}_{1|1}\in L(1,1)$, $\mathfrak{b}_{n-1|n-1}\in L(n-1,n-1)$, and $\lambda\in\mathfrak{h}^{*}$ with restrictions $\lambda|_{1|1}$ and $\lambda|_{n-1|n-1}$ as in Remark 5. Then there is an isomorphism
$\mathrm{Ind}_{\mathfrak{p}_{IJ}}^{\mathfrak{g}}\mathrm{Infl}_{\mathfrak{l}_{IJ}}^{\mathfrak{p}_{IJ}}\Bigl(M^{\mathfrak{b}_{1|1}}\bigl(\lambda|_{1|1}\bigr)\ \boxtimes\ M^{\mathfrak{b}_{n-1|n-1}}\bigl(\lambda|_{n-1|n-1}\bigr)\Bigr)\ \cong\ M^{\mathfrak{b}_{1|1}\star\mathfrak{b}_{n-1|n-1}}(\lambda).$
@@
$\operatorname{Ind}_{\mathfrak{p}_{IJ}}^{\mathfrak{gl}(n|n)}\operatorname{Infl}_{\mathfrak{l}_{IJ}}^{\mathfrak{p}_{IJ}}\Bigl(L^{()}(\lambda|_{1|1})\boxtimes\operatorname{BG}_{n-1|n-1}(\lambda|_{n-1|n-1})\Bigr)\ \cong\ \operatorname{BG}_{n|n}(\lambda).$
@@
To make this explicit, define a subsuperalgebra
@@
$\operatorname{DS}_{e_{1,n+1}}\Bigl(M^{()\star\mathfrak{b}_{n-1|n-1}}(\lambda)\Bigr)\ \cong\ \begin{cases}M^{\mathfrak{b}_{n-1|n-1}}\bigl(\lambda|_{n-1|n-1}\bigr)\oplus\Pi\,M^{\mathfrak{b}_{n-1|n-1}}\bigl(\lambda|_{n-1|n-1}\bigr),&\lambda|_{1|1}\ \text{atypical},\\ 0,&\lambda|_{1|1}\ \text{typical}.\end{cases}$
@@
$\operatorname{DS}_{e_{1,n+1}}\Bigl(M^{\mathfrak{b}_{\mathbf{o}}}(\lambda)\Bigr)\ \cong\ \begin{cases}M^{\mathfrak{b}_{\mathbf{o}}}\bigl(\lambda|_{n-1|n-1}\bigr)\oplus\Pi\,M^{\mathfrak{b}_{\mathbf{o}}}\bigl(\lambda|_{n-1|n-1}\bigr),&\lambda|_{1|1}\ \text{atypical},\\ 0,&\lambda|_{1|1}\ \text{typical}.\end{cases}$
@@
$\operatorname{DS}_{e_{1,n+1}}\Bigl(M^{()\star\mathfrak{b}_{n-1|n-1}}(\lambda)\Bigr)$ $\cong\operatorname{DS}_{e_{1,n+1}}\operatorname{Ind}_{\mathfrak{p}_{IJ}}^{\mathfrak{g}}\operatorname{Infl}_{\mathfrak{l}_{IJ}}^{\mathfrak{p}_{IJ}}\Bigl(M^{\mathfrak{b}_{1|1}}\bigl(\lambda|_{1|1}\bigr)\boxtimes M^{\mathfrak{b}_{n-1|n-1}}\bigl(\lambda|_{n-1|n-1}\bigr)\Bigr)$
$\cong\operatorname{DS}^{\mathfrak{l}_{IJ}}_{e_{1,n+1}}\Bigl(M^{\mathfrak{b}_{1|1}}\bigl(\lambda|_{1|1}\bigr)\boxtimes M^{\mathfrak{b}_{n-1|n-1}}\bigl(\lambda|_{n-1|n-1}\bigr)\Bigr)$
$\cong\Bigl(\operatorname{DS}_{e_{1,n+1}}M^{\mathfrak{b}_{1|1}}\bigl(\lambda|_{1|1}\bigr)\Bigr)\boxtimes M^{\mathfrak{b}_{n-1|n-1}}\bigl(\lambda|_{n-1|n-1}\bigr).$
Here the first isomorphism follows from Lemma 5, the second from Theorem 5.1, and the third from Example 4. Now apply Example 4 to $\operatorname{DS}_{e_{1,n+1}}M^{\mathfrak{b}_{1|1}}\bigl(\lambda|_{1|1}\bigr)$ to obtain the stated cases. $\square$
@@
$\operatorname{DS}_{e_{1,n+1}}\bigl(\operatorname{BG}_{n|n}(\lambda)\bigr)\ \cong\ \Pi^{\operatorname{par}(\lambda)}\,\operatorname{BG}_{n-1|n-1}\bigl(\lambda|_{n-1|n-1}\bigr).$
@@
$0\to 0\to\Pi\,\Pi^{\operatorname{par}(\lambda|_{1|1})}L^{()}(b\mid b)\to L^{()}(b\mid b)\oplus\Pi L^{()}(b\mid b)\to\Pi^{\operatorname{par}(\lambda|_{1|1})}L^{()}(b\mid b)\to 0\to 0.$
@@
$0\to 0\to\Pi^{\operatorname{par}(\lambda|_{1|1})}L^{()}(b-1\mid b-1)\to\Pi^{\operatorname{par}(\lambda|_{1|1})}M^{()}(b\mid b)\to\Pi^{\operatorname{par}(\lambda|_{1|1})}L^{()}(b\mid b)\to 0\to 0.$
@@
$0\to 0\to\Pi\,\Pi^{\operatorname{par}(\lambda|_{1|1})}M^{()}(b\mid b)\to M^{()}(b\mid b)\oplus\Pi M^{()}(b\mid b)\to\Pi^{\operatorname{par}(\lambda|_{1|1})}M^{()}(b\mid b)\to 0\to 0.$
@@
$0\to 0\to\Pi\,\Pi^{\operatorname{par}(\lambda|_{1|1})}L^{()}(b\mid b)\to 0\to\Pi^{\operatorname{par}(\lambda|_{1|1})}L^{()}(b\mid b)\to 0\to 0.$
@@
$0\to 0\to\Pi^{\operatorname{par}(\lambda|_{1|1})}L^{()}(b+1\mid b+1)\to\Pi^{\operatorname{par}(\lambda|_{1|1})}M^{(1)}(b+1\mid b+1)\to\Pi^{\operatorname{par}(\lambda|_{1|1})}L^{()}(b\mid b)\to 0\to 0.$
@@
$0\to 0\to\Pi\,\Pi^{\operatorname{par}(\lambda|_{1|1})}M^{()}(b\mid b)\to 0\to\Pi^{\operatorname{par}(\lambda|_{1|1})}M^{()}(b\mid b)\to 0\to 0.$

### v2

# Duflo-Serganova functors and Brundan-Goodwin’s parabolic inductions

@@
Duflo–Serganova functors play an important role in the representation theory of Lie superalgebra. While it is desirable to understand the images of modules under DS, little is known beyond finite dimensional representations. For general linear Lie superalgebras, Brundan–Goodwin observed that certain parabolic induction functors are compatible with the principal Whittaker coinvariant functor $H_{0}$.
In this paper we show that, for certain rank-one DS functors, we explicitly compute the image of these parabolically induced modules under DS explicitly. This class of module includes many infinite dimensional representations such as $\mathfrak{b}$-Verma supermodules (attached to a suitable Borel subal gebra $\mathfrak{b}$) and a family of modules whose $H_{0}$-images are tensor products of irreducible evaluation modules for the super Yangian.
@@ 1 Introduction
Highest weight representations are basic objects in the representation theory of semisimple Lie algebras. In the study of these representations, endofunctors attached to simple roots—such as translation functors—have played a crucial role [17]. Representation theory of basic classical Lie superalgebras, which includes the general linear Lie superalgebra $\mathfrak{gl}(m|n)$ [6], naturally extends the classical theory and exhibits new phenomena absent in the Lie algebra case, such as the data of Borel subalgebras.
@@
One of the examples of new constructions appearing in the Lie superalgebra setting is the Duflo–Serganova functor [10] (DS for short), which is a “good” functor attached to an odd root. For $\mathfrak{gl}(m|n)$ it maps representations to those of a smaller Lie superalgebra $\mathfrak{gl}(m-r|n-r)$. This functor is symmetric monoidal and has implications to the theory of finite-dimensional representations; see e.g., [12]. In addition, Heidersdorf–Weissauer [14] determine the images of finite-dimensional irreducible $\mathfrak{gl}(n|n)$-modules under DS in terms of Khovanov arc diagrams.
@@
In contrast, much less is known for behavior of infinite-dimensional representations under DS. In [9], Coulembier and Serganova studied when Verma modules annihilated by DS, and apply this to homological-algebraic questions (for its extension to arbitrary Borel subalgebras, see [15]). Hoyt–Penkov–Serganova [16] show that DS preserve the category of highest weight representations and formulate further conjectures on their behavior. Taken together, these developments suggest that it is natural to investigate the behavior of DS on general highest weight representations. To the best of the author’s knowledge, little is known about this. The aim of this paper is to spell out explicit formula for a class of infinite-dimensional highest weight representations of $\mathfrak{gl}(n|n)$ with respect to a rank-one Duflo-Serganova functor to improve the situation.
@@
Hypercubic decomposition of Verma modules. We first study the $\mathfrak{b}$-Verma modules (Verma module with respect to a Borel subalgebra $\mathfrak{b}$ of $\mathfrak{gl}(n|n)$). In view of the fact that DS is middle exact, it is natural to decompose $\mathfrak{b}$-Verma modules along the directions of odd $\mathfrak{b}$-simple roots. This idea leads to a realization of an abelian category equivalent to the direct sum of the category of highest weight representations of $\mathfrak{gl}(1|1))$ inside the category $\mathcal{O}_{\mathfrak{gl}(n|n)}$ of highest weight representations of $\mathfrak{gl}(n|n)$ [15] (see Cheng–Lam–Wang [8] and Serganova [19] for related works). Pursuing this direction leads to a family of hypercube Borels, for which some of the associated $\mathfrak{b}$-Verma modules can be realized inside an abelian subcategory equivalent to the maximal atypical block of $\mathcal{O}_{\mathfrak{gl}(1|1)^{\oplus n}}$.
@@
For example, for $\mathfrak{gl}(3|3)$ one can visualize the collection of Borel subalgebras with fixed even part as a finite graph whose vertices are Borels and whose edges correspond to odd reflections. In this picture there is a distinguished “cube” in the middle, and the family of hypecube Borels refers to the Borel subalgebras corresponding to the vertices of this central cube. Note that here the uppertrianglar Borel subalgebra corresponds to $()$.
()()(1)(1)(21)(21)(22)($^{2}$$)(1^{2}$$)(1^{3}$$1^{2})(2)(2$)(3)(3)(31)(31)(33)($^{3}$$)(3^{2}2$$)(312)(31^{2}$$^{2})(23)(2^{3}$$)(2^{2}1$$)(3^{2}1$$)(3^{2}$)$(32)$
@@
Parabolic induction and Whittaker coinvariant functor. Decomposition of Verma modules for hypercube Borels respects the internal structure of parabolic induction functor of Brundan–Goodwin [4], which sends $\mathfrak{gl}(1|1)^{\oplus n}$-modules to $\mathfrak{gl}(n|n)$-modules. They are motivated by studying the principal Whittaker coinvariant functor $H_{0}$, which sends objects of category $\mathcal{O}$ to modules over the principal $W$-superalgebra, which is nontrivial contrary to the Lie algbera setting. Generalizing [5], Brown–Brundan–Goodwin [3] show that the principal $W$-superalgebra can be realized as a quotient of the super Yangian $Y(\mathfrak{gl}(1|1))$.
Main Restult.
We conjecture the following description of the effect of DS functors on Verma modules.
@@
We prove this conjecture in several settings. In particular, in Theorem 5.7 we verify it for hypercube Borels with $e_{\alpha}=e_{1,n+1}$, and by direct computations, in Theorem 6.6 we verify it for $\mathfrak{gl}(2|2)$ for an arbitrary Borel subalgebra and an arbitrary $\mathfrak{b}$-simple odd root.
In a sense, our conjecture say computations for $\mathfrak{b}$-Verma modules are independent of the choice of Borel subalgebras. As a special feature of hypercube Borels, we prove the following result.
@@
$\operatorname{DS}_{e_{1,n+1}}\operatorname{BG}_{n|n}(\lambda)\ \cong\ \operatorname{BG}_{n-1|n-1}\bigl(\operatorname{pr}_{\alpha}(\lambda)\bigr).$
Moreover $H_{0}\operatorname{BG}_{n-1|n-1}\bigl(\operatorname{pr}_{\alpha}(\lambda)\bigr)$ is also one-dimensional.
Remark. The main objects of study in this paper, namely the modules $\operatorname{BG}(\lambda)$, can be realized as images of homomorphisms between Verma modules attached to Borel subalgebras lying at opposite vertices of the central hypercube.
Finally, we note that, from the perspective of Nichols algebras of diagonal type and their classification via Weyl groupoids (a.k.a. generalized root systems), dependence of highest weight structures with respect to the choice of Borel subalgebras are of independent interest (see [2, 1, 13]. We believe that the results of this paper (Theorems 3.13 and 5.9) provide further evidence for the significance of such perspective.
@@ 1.1 Acknowledgments
We define the full subcategory of $\mathcal{O}$ consisting of modules whose restriction admit an even Verma flag, and denote it by $\mathcal{F}\!\Delta_{\bar{0}}\subset\mathcal{O}.$
@@
    $L^{()}(\lambda)$ injectively maps to some $M\in\mathcal{F}\!\Delta_{\bar{0}}.$
@@ 3 Brundan-Goodwin functors
Under the principal Whittaker coinvariants functor $H_{0}$, the module $\operatorname{BG}(\lambda)$ is sent to the corresponding tensor product of irreducible evaluation module for the (super) Yangian. Moreover,
@@
    $\operatorname{BG}(\lambda)\ \cong\ \operatorname{Im}\!\left(M^{\mathfrak{b}_{\mathbf{o}}}(\lambda)\xrightarrow{\ \neq 0\ }M^{\mathfrak{b}_{\mathbf{i}}}(\lambda-ber)\right).$
@@ 4 Duflo-Serganova functors
For $x\in X$ , $M\in\mathfrak{g}\text{-sMod}$, we define a supervector space $DS_{x}M$ by:
@@
naturally inherits the structure of a Lie superalgebra. Consequently, there is a natural symmetric monoidal $k$-linear functor
@@ 5 Main results
Let $\mathfrak{b}_{1|1}=(\beta_{n})\in L(1,1)$ and $\mathfrak{b}_{n-1|n-1}=(\beta_{1},\dots,\beta_{n-1})$. Define
@@
For $\lambda\in\Lambda_{n|n}$ we write $\operatorname{pr}_{J}(\lambda)\in\Lambda_{n-1|n-1}$ and $\operatorname{pr}_{I}(\lambda)\in\Lambda_{1|1}$ for the restrictions corresponding to the Levi embedding
@@
$\operatorname{pr}_{J}(\lambda)\in\Lambda^{\operatorname{BG}}_{n-1|n-1}\qquad\text{(resp.\ }\operatorname{pr}_{J}(\lambda)\in\Lambda^{\operatorname{maBG}}_{n-1|n-1}\text{)}.$
@@
Let $\mathfrak{b}_{1|1}\in L(1,1)$, $\mathfrak{b}_{n-1|n-1}\in L(n-1,n-1)$, and $\lambda\in\mathfrak{h}^{*}$ with restrictions $\operatorname{pr}_{I}(\lambda)$ and $\operatorname{pr}_{J}(\lambda)$ as in Remark 5. Then there is an isomorphism
$\mathrm{Ind}_{\mathfrak{p}_{IJ}}^{\mathfrak{g}}\mathrm{Infl}_{\mathfrak{l}_{IJ}}^{\mathfrak{p}_{IJ}}\Bigl(M^{\mathfrak{b}_{1|1}}\bigl(\operatorname{pr}_{I}(\lambda)\bigr)\ \boxtimes\ M^{\mathfrak{b}_{n-1|n-1}}\bigl(\operatorname{pr}_{J}(\lambda)\bigr)\Bigr)\ \cong\ M^{\mathfrak{b}_{1|1}\star\mathfrak{b}_{n-1|n-1}}(\lambda).$
@@
$\operatorname{Ind}_{\mathfrak{p}_{IJ}}^{\mathfrak{gl}(n|n)}\operatorname{Infl}_{\mathfrak{l}_{IJ}}^{\mathfrak{p}_{IJ}}\Bigl(L^{()}(\operatorname{pr}_{I}(\lambda))\boxtimes\operatorname{BG}_{n-1|n-1}(\operatorname{pr}_{J}(\lambda))\Bigr)\ \cong\ \operatorname{BG}_{n|n}(\lambda).$
@@
Define a subsuperalgebra
@@
$\operatorname{DS}_{e_{1,n+1}}\Bigl(M^{()\star\mathfrak{b}_{n-1|n-1}}(\lambda)\Bigr)\ \cong\ \begin{cases}M^{\mathfrak{b}_{n-1|n-1}}\bigl(\operatorname{pr}_{J}(\lambda)\bigr)\oplus\Pi\,M^{\mathfrak{b}_{n-1|n-1}}\bigl(\operatorname{pr}_{J}(\lambda)\bigr),&\operatorname{pr}_{I}(\lambda)\ \text{atypical},\\ 0,&\operatorname{pr}_{I}(\lambda)\ \text{typical}.\end{cases}$
@@
$\operatorname{DS}_{e_{1,n+1}}\Bigl(M^{\mathfrak{b}_{\mathbf{o}}}(\lambda)\Bigr)\ \cong\ \begin{cases}M^{\mathfrak{b}_{\mathbf{o}}}\bigl(\operatorname{pr}_{J}(\lambda)\bigr)\oplus\Pi\,M^{\mathfrak{b}_{\mathbf{o}}}\bigl(\operatorname{pr}_{J}(\lambda)\bigr),&\operatorname{pr}_{I}(\lambda)\ \text{atypical},\\ 0,&\operatorname{pr}_{I}(\lambda)\ \text{typical}.\end{cases}$
@@
$\operatorname{DS}_{e_{1,n+1}}\Bigl(M^{()\star\mathfrak{b}_{n-1|n-1}}(\lambda)\Bigr)$ $\cong\operatorname{DS}_{e_{1,n+1}}\operatorname{Ind}_{\mathfrak{p}_{IJ}}^{\mathfrak{g}}\operatorname{Infl}_{\mathfrak{l}_{IJ}}^{\mathfrak{p}_{IJ}}\Bigl(M^{\mathfrak{b}_{1|1}}\bigl(\operatorname{pr}_{I}(\lambda)\bigr)\boxtimes M^{\mathfrak{b}_{n-1|n-1}}\bigl(\operatorname{pr}_{J}(\lambda)\bigr)\Bigr)$
$\cong\operatorname{DS}^{\mathfrak{l}_{IJ}}_{e_{1,n+1}}\Bigl(M^{\mathfrak{b}_{1|1}}\bigl(\operatorname{pr}_{I}(\lambda)\bigr)\boxtimes M^{\mathfrak{b}_{n-1|n-1}}\bigl(\operatorname{pr}_{J}(\lambda)\bigr)\Bigr)$
$\cong\Bigl(\operatorname{DS}_{e_{1,n+1}}M^{\mathfrak{b}_{1|1}}\bigl(\operatorname{pr}_{I}(\lambda)\bigr)\Bigr)\boxtimes M^{\mathfrak{b}_{n-1|n-1}}\bigl(\operatorname{pr}_{J}(\lambda)\bigr).$
Here the first isomorphism follows from Lemma 5, the second from Theorem 5.1, and the third from Example 4. Now apply Example 4 to $\operatorname{DS}_{e_{1,n+1}}M^{\mathfrak{b}_{1|1}}\bigl(\operatorname{pr}_{I}(\lambda)\bigr)$ to obtain the stated cases. $\square$
@@
$\operatorname{DS}_{e_{1,n+1}}\bigl(\operatorname{BG}_{n|n}(\lambda)\bigr)\ \cong\ \Pi^{\operatorname{par}(\operatorname{pr}_{I}(\lambda))}\,\operatorname{BG}_{n-1|n-1}\bigl(\operatorname{pr}_{J}(\lambda)\bigr).$
@@
$0\to 0\to\Pi\,\Pi^{\operatorname{par}(a|a)}L^{()}(b\mid b)\to L^{()}(b\mid b)\oplus\Pi L^{()}(b\mid b)\to\Pi^{\operatorname{par}(a|a)}L^{()}(b\mid b)\to 0\to 0.$
@@
$0\to 0\to\Pi^{\operatorname{par}(a|a)}L^{()}(b-1\mid b-1)\to\Pi^{\operatorname{par}(a|a)}M^{()}(b\mid b)\to\Pi^{\operatorname{par}(a|a)}L^{()}(b\mid b)\to 0\to 0.$
@@
$0\to 0\to\Pi\,\Pi^{\operatorname{par}(a|a)}M^{()}(b\mid b)\to M^{()}(b\mid b)\oplus\Pi M^{()}(b\mid b)\to\Pi^{\operatorname{par}(a|a)}M^{()}(b\mid b)\to 0\to 0.$
@@
$0\to\Pi\,\Pi^{\operatorname{par}(a|a)}L^{()}(b\mid b)\to\Pi\,\Pi^{\operatorname{par}(a|a)}L^{()}(b\mid b)\to 0\to\Pi^{\operatorname{par}(a|a)}L^{()}(b\mid b)\to\Pi^{\operatorname{par}(a|a)}L^{()}(b\mid b)\to 0.$
@@
$0\to 0\to\Pi^{\operatorname{par}(a|a)}L^{()}(b+1\mid b+1)\to\Pi^{\operatorname{par}(a|a)}M^{(1)}(b+1\mid b+1)\to\Pi^{\operatorname{par}(a|a)}L^{()}(b\mid b)\to 0\to 0.$
@@
$0\to\Pi\,\Pi^{\operatorname{par}(a|a)}M^{()}(b\mid b)\to\Pi\,\Pi^{\operatorname{par}(a|a)}M^{()}(b\mid b)\to 0\to\Pi^{\operatorname{par}(a|a)}M^{()}(b\mid b)\to\Pi^{\operatorname{par}(a|a)}M^{()}(b\mid b)\to 0.$